# Supertranslation in the bulk for generic spacetime

### v1

@@ Abstract
Supertranslations are usually defined as asymptotic symmetries associated with spacetime boundaries, such as null infinity and black hole horizons. In this Letter, we show that supertranslations admit a natural, coordinate-independent extension into the bulk of spacetime, realized as transitions between families of null hypersurfaces. This construction applies to generic spacetimes in arbitrary dimensions and unifies the realizations of supertranslations at null infinity and black hole horizons. The associated symmetry algebra can be realized by light-ray operators defined on the null hypersurface. Within this framework, the gravitational wave memory effect corresponds to a shift of null hypersurfaces in the bulk. As explicit examples, we compute bulk supertranslations in Minkowski spacetime in arbitrary dimensions and in four-dimensional Schwarzschild spacetime, where we uncover a novel curvature-induced memory effect with observable consequences for light propagation.
Introduction. Supertranslation is a very surprising result from the investigation on asymptotic structure of the spacetime at null infinity Bondi et al. (1962); Sachs (1962a, b), where the spacetime symmetry is enhanced from the Poincaré group to the Bondi-Metzner-Sachs (BMS) group. The BMS group is the semidirect product of the Lorentz group and an infinite-dimensional Abelian group, the supertranslation that generalizes the ordinary translations. Although supertranslation symmetry was discovered long ago, its physical implications have only been appreciated recently Strominger (2018). The major breakthrough was made by Strominger for the discovery of the BMS invariance of gravitational scattering Strominger (2014). Supertranslations lie at the heart of a fascinating triangular equivalence Strominger (2018). It reveals the symmetry origin of the Weinberg’s soft graviton theorem at the quantum level He et al. (2015) and characterizes the vacua transition associated with the gravitational memory effect at the classical level Strominger and Zhiboedov (2016). Further developments from various perspectives can be found, e.g., in Bousso and Porrati (2017); Sousa et al. (2018); Choi and Akhoury (2018); Córdova and Shao (2018); Donnay et al. (2020); Chen et al. (2021); Fuentealba et al. (2021); Chakraborty et al. (2022); Fuentealba et al. (2022); Veneziano and Vilkovisky (2022); Javadinezhad and Porrati (2023); Fuentealba et al. (2023); Henneaux (2024); Javadinezhad and Porrati (2024); Elkhidir et al. (2025); Henneaux (2025).
@@
Motivated by the intrinsic connection between supertranslations and gravitational memory, we propose a bulk memory effect corresponding to the transition of null geodesics from one hypersurface to another. Thus, supertranslations are not merely asymptotic but define a genuine bulk symmetry realized by null hypersurface flows with observable memory effects. As concrete demonstrations, we compute supertranslations for Minkowski spacetime in four dimensions, which recovers the bulk extension of supertranslations in Minkowski spacetime obtained in Compère and Long (2016a). The computations are then extended to Minkowski spacetime in arbitrary dimensions.
A novel curvature-dependent memory effect is uncovered from the supertranslation in four-dimensional Schwarzschild spacetime, where elliptic integrals are involved with branch points that correspond to the turning points of null geodesics on the hypersurface. Specifically, light rays initially placed on a hypersurface without turning points can be shifted to a new hypersurface with turning points by gravitational wave memory, which allows them to return from the bulk. This type of memory effect occurs only in curved spacetime, i.e., a black hole memory Donnay et al. (2018); Rahman and Wald (2020). In flat spacetime, light rays remain straight, so bulk memory (supertranslations) can distort their trajectories but never generate turning points.
@@
Residual gauge transformation and supertranslation. In this section, we first examine the physical implications of residual gauge transformations within the NU gauge and identify the supertranslations in this context. We then generalize these definitions to a generic coordinate system. Suppose that two NU coordinate systems $(u,r,x^{a})$ and $(\bar{u},\bar{r},\bar{x}^{a})$ are related by $\bar{x}^{\mu}=\bar{x}^{\mu}(x^{\alpha})$. A key feature of the NU gauge is the choice of a null hypersurface labeled by a single coordinate. If the spacetime can be described in both NU coordinate systems, then the hypersurfaces u=u=const and $\bar{u}=$const are both null. Consequently, the full residual gauge transformations of the NU gauge are characterized by different families of null hypersurfaces. In the NU coordinates $(u,r,x^{a})$, a residual gauge transformation is determined by a scalar potential $\bar{u}$, with the condition that $\bar{\ell}=\text{d}\bar{u}$ is a null vector.
@@
  • Applying a near-horizon series expansion, the supertranslation reduces to the near horizon supertranslations identified in Donnay et al. (2016a, b). This generic definition of supertranslation thus unifies the null infinity and near-horizon symmetries, providing a natural framework for matching conserved quantities at different boundaries, see, e.g., relevant investigations in Hawking et al. (2016); Grumiller et al. (2020); Ruzziconi and Zwikel (2025a); Agrawal et al. (2025); Ruzziconi and Zwikel (2025b).
@@
One may initially select a family of null hypersurfaces on which the null geodesics exhibit no turning points. Gravitational waves with memory can then shift the system to a different null hypersurface where the light rays develop turning points. As a result, it becomes, in principle, possible to receive light rays returning from the bulk of spacetime. This constitutes a new type of gravitational memory effect that arises purely from spacetime curvature.
@@
Figure 2: The black region represents the black hole. The black wavy lines indicate gravitational waves with memory, while the blue lines show the null geodesic before and after the memory effect.
@@
We determine the full residual gauge transformations associated with the choice of null hypersurfaces, which consist of supertranslations and a Weyl-type rescaling of null vectors. In the present work, we focus on the supertranslation sector. It is of independent interest to investigate the rescaling factor $\Phi$, which account for various extensions of the BMS symmetry. Another important direction is to clarify the relation between the bulk supertranslations constructed here and those arising in other contexts, such as scattering amplitudes Elkhidir et al. (2025) and holography Marjieh et al. (2022). A particular notable feature of our construction is that bulk supertranslations are realized as characteristic flows of boundary supertranslations. In the context of AdS/CFT, the gravity dual of a deformation by the composite operator $T\bar{T}$ Smirnov and Zamolodchikov (2017); Cavaglià et al. (2016) (also Jiang (2021); He et al. (2025) for reviews) corresponds to moving the boundary to a finite radial cutoff McGough et al. (2018); Kraus et al. (2018); Hartman et al. (2019); Guica and Monten (2021); Jafari et al. (2020); He et al. (2023, 2024). The associated $T\bar{T}$ flow has been identified as a characteristic flow Hou (2023). These parallels suggest that our method of flowing asymptotic symmetries into the bulk is naturally compatible with the AdS/CFT perspective and merits further investigation.
Acknowledgments. This work is supported in part by the National Natural Science Foundation of China (NSFC) under Grants No. 12475059 and No. 11935009, and by Tianjin University Self-Innovation Fund Extreme Basic Research Project Grant No. 2025XJ21-0007.

### v2

@@ Abstract
Supertranslations are usually defined as asymptotic symmetries associated with spacetime boundaries, such as null infinity and black hole horizons. In this Letter, we show that supertranslations admit a natural, coordinate-independent extension into the bulk of spacetime, realized as transitions between families of null hypersurfaces. This construction applies to generic spacetimes in arbitrary dimensions and unifies the realizations of supertranslations at null infinity and black hole horizons. The bulk supertranslation is connected to boundary supertranslation by characteristic flows. The associated symmetry algebra can be realized by light-ray operators defined on the null hypersurface. Within this framework, the gravitational wave memory effect corresponds to a shift of null hypersurfaces in the bulk. As explicit examples, we compute bulk supertranslations in Minkowski spacetime in arbitrary dimensions and in four-dimensional Schwarzschild spacetime, where we uncover a novel curvature-induced memory effect with observable consequences for light propagation.
Introduction. Supertranslation is a very surprising result from the investigation on asymptotic structure of the spacetime at null infinity Bondi et al. (1962); Sachs (1962a, b), where the spacetime symmetry is enhanced from the Poincaré group to the Bondi-Metzner-Sachs (BMS) group. The BMS group is the semidirect product of the Lorentz group and an infinite-dimensional Abelian group, the supertranslation that generalizes the ordinary translations. Although supertranslation symmetry was discovered long ago, its physical implications have only been appreciated recently Strominger (2018). The major breakthrough was made by Strominger for the discovery of the BMS invariance of gravitational scattering Strominger (2014). Supertranslations lie at the heart of a fascinating triangular equivalence Strominger (2018). It reveals the symmetry origin of the Weinberg’s soft graviton theorem at the quantum level He et al. (2015) and characterizes the vacua transition associated with the gravitational memory effect at the classical level Strominger and Zhiboedov (2016). Further developments from various perspectives can be found, e.g., in Carlip (2018); Bousso and Porrati (2017); Sousa et al. (2018); Choi and Akhoury (2018); Córdova and Shao (2018); Donnay et al. (2020); Chen et al. (2021); Fuentealba et al. (2021); Chakraborty et al. (2022); Fuentealba et al. (2022); Veneziano and Vilkovisky (2022); Javadinezhad and Porrati (2023); Fuentealba et al. (2023); Henneaux (2024); Javadinezhad and Porrati (2024); Elkhidir et al. (2025); De Luca et al. (2025a, b); Henneaux (2025).
@@
Motivated by the intrinsic connection between supertranslations and gravitational memory, we propose a bulk memory effect corresponding to the transition of null geodesics from one hypersurface to another. Thus, supertranslations are not merely asymptotic but define a genuine bulk symmetry with observable memory effects Goncharov et al. (2024). As concrete demonstrations, we compute supertranslations for Minkowski spacetime in four dimensions, which recovers the bulk extension of supertranslations in Minkowski spacetime obtained in Compère and Long (2016a). The computations are then extended to Minkowski spacetime in arbitrary dimensions.
A novel curvature-dependent memory effect is uncovered from the supertranslation in four-dimensional Schwarzschild spacetime, where elliptic integrals are involved with branch points that correspond to the turning points of null geodesics on the hypersurface. Specifically, light rays initially placed on a hypersurface without turning points can be shifted to a new hypersurface with turning points by gravitational wave memory, which allows them to return from the bulk. This type of memory effect occurs only in curved spacetime flat , namely black hole memory Donnay et al. (2018); Rahman and Wald (2020), and can be directly determined within black hole perturbation theory, such as gravitational self-force theory Elhashash and Nichols (2025); Cunningham et al. (2025).
@@
Residual gauge transformation and supertranslation. In this section, we first examine the physical implications of residual gauge transformations within the NU gauge and identify the supertranslations in this context. We then generalize these definitions to a generic coordinate system. Suppose that two NU coordinate systems $(u,r,x^{a})$ and $(\bar{u},\bar{r},\bar{x}^{a})$ are related by $\bar{x}^{\mu}=\bar{x}^{\mu}(x^{\alpha})$. A key feature of the NU gauge is the choice of a null hypersurface labeled by a single coordinate. If the spacetime can be described in both NU coordinate systems, then the hypersurfaces u=u=const and $\bar{u}=$const are both null. Consequently, the full residual gauge transformations of the NU gauge are characterized by different families of null hypersurfaces hypersurface . In the NU coordinates $(u,r,x^{a})$, a residual gauge transformation is determined by a scalar potential $\bar{u}$, with the condition that $\bar{\ell}=\text{d}\bar{u}$ is a null vector.
@@
  • Applying a near-horizon series expansion, the supertranslation reduces to the near horizon supertranslations identified in Donnay et al. (2016a, b), and also Afshar et al. (2016); Shi and Mei (2017); Akhmedov and Godazgar (2017); Chandrasekaran et al. (2018); Haco et al. (2018); Grumiller et al. (2020a); Adami et al. (2020c, a); Chen et al. (2020); Adami et al. (2020b); Chandrasekaran et al. (2022); Adami et al. (2021) for other realizations. This generic definition of supertranslation thus unifies the null infinity and near-horizon symmetries, providing a natural framework for matching conserved quantities at different boundaries, see, e.g., relevant investigations in Hawking et al. (2016); Grumiller et al. (2020); Ruzziconi and Zwikel (2025a); Agrawal et al. (2025); Ruzziconi and Zwikel (2025b).
@@
One may initially select a family of null hypersurfaces on which the null geodesics exhibit no turning points. Gravitational waves with memory can then shift the system to a different null hypersurface where the light rays develop turning points. As a result, it becomes, in principle, possible to receive light rays returning from the bulk of spacetime. This constitutes a new type of gravitational memory effect that arises purely from spacetime curvature. Such memory effect can be directly computed within black hole perturbation theory, without the need to refer to a flat spacetime background or to null infinity.
@@
Figure 2: The black region represents the black hole. The black wavy lines indicate gravitational waves with memory, while the blue lines show the null geodesic before and after the memory effect. The memory effect can be observed from the propagation of one light ray.
@@
We determine the full residual gauge transformations associated with the choice of null hypersurfaces, which consist of supertranslations and a Weyl-type rescaling of null vectors. In the present work, we focus on the supertranslation sector. It is of independent interest to investigate the rescaling factor $\Phi$, which account for various extensions of the BMS symmetry. Another important direction is to clarify the relation between the bulk supertranslations constructed here and those arising in other contexts, such as scattering amplitudes Elkhidir et al. (2025) and holography Marjieh et al. (2022). A particular notable feature of our construction is that bulk supertranslations are realized as characteristic flows of boundary supertranslations. In the context of AdS/CFT, the gravity dual of a deformation by the composite operator $T\bar{T}$ Smirnov and Zamolodchikov (2017); Cavaglià et al. (2016) (also Jiang (2021); He et al. (2025) for reviews) corresponds to moving the boundary to a finite radial cutoff McGough et al. (2018); Kraus et al. (2018); Hartman et al. (2019); Guica and Monten (2021); Jafari et al. (2020); He et al. (2023, 2024); moving . The associated $T\bar{T}$ flow has been identified as a characteristic flow Hou (2023). These parallels suggest that our method of flowing asymptotic symmetries into the bulk is naturally compatible with the AdS/CFT perspective and merits further investigation.
Acknowledgments. The author would like to thank Hamed Adami, Jia Du, Shaoqi Hou, Delong Kong, Shahin Sheikh-Jabbari, Vahid Taghiloo, Yu Tian, and Hongbao Zhang for useful discussions. This work is supported in part by the National Natural Science Foundation of China (NSFC) under Grants No. 12475059 and No. 11935009, and by Tianjin University Self-Innovation Fund Extreme Basic Research Project Grant No. 2025XJ21-0007.